OpenStudy (anonymous):

4-2(v-6)=-8 Please? Algebra help?

OpenStudy (anonymous):

Please Hurry :(

OpenStudy (anonymous):

24-2v=0 -2(v-8)=-8 16-2v=-8 V=12

OpenStudy (anonymous):

Ohhh... I get it now Thank you!!!